OpenStudy (anonymous):

A candle is 17 in. tall after burning for 3 hours. After 5 hours, it is 15in tall. Write a linear equation to model the relationship between height h of the candle and time t. Predict how tall the candle will be after burning 8 hours. **WILL GIVE MEDAL TO #1 HELPER**

OpenStudy (anonymous):

@zepdrix I know you've done this one before! :D

zepdrix (zepdrix):

Oh man the stupid candle question again? +_+ haha

zepdrix (zepdrix):

sec, brb

zepdrix (zepdrix):

ok ok ok ok ok ok something about candle?

OpenStudy (anonymous):

Yes. im confused on the formula? if there even is one.

zepdrix (zepdrix):

So they told us that the relationship can be modeled by a `linear equation`. A linear equation is of the form:\[\Large y=mx+b\]Where m is the slope of the line, and b is the y-intercept.

zepdrix (zepdrix):

Since we're dealing with height and time, instead of the normal x and y, let's plug in different variables. Hopefully this won't get too confusing for you :)\[\Large h=mt+b\]

zepdrix (zepdrix):

The information they gave us about the height at specific times we can think of as coordinate pairs.\[\Large (t_1,\;h_1)\quad=\quad (3,\;17)\]

zepdrix (zepdrix):

\[\Large (t_2,\;h_2)\quad=\quad (5,\;15)\]

zepdrix (zepdrix):

We can use these coordinate pairs to find the slope of our line. Remember how to find slope? :)

OpenStudy (anonymous):

Slope isn't y = mx + b, is it?

OpenStudy (anonymous):

So slope is m, m = (y_f - y_i) / (x_f - x_i)

zepdrix (zepdrix):

slope is\[\Large m=\frac{y_2-y_1}{x_2-x_1}\]

OpenStudy (anonymous):

Oh, right right right.

zepdrix (zepdrix):

Change in y, over change in x :D we're using h and t though, so our setup will look like:\[\Large m\quad=\quad\frac{h_2-h_1}{t_2-t_1}\]

OpenStudy (anonymous):

Alright, Im on top of things. What next?

OpenStudy (anonymous):

Do I just plug-in now?

zepdrix (zepdrix):

Plug in the m, yes.

zepdrix (zepdrix):

What'd we end up with? :)

OpenStudy (anonymous):

One sec.

OpenStudy (anonymous):

uuuh...I got 7/5 (1.4)? 17-3/15-5

zepdrix (zepdrix):

Hmm your setup looks a little strange.\[\Large h_2=17, \qquad h_1=15\]So the top of our slope formula should look like 17-15

OpenStudy (anonymous):

Oh, that makes more sense. okay one sec again

OpenStudy (anonymous):

-1?

zepdrix (zepdrix):

cool, sounds right :)

OpenStudy (anonymous):

WOO! :D

zepdrix (zepdrix):

So plugging in our m, we currently have:\[\Large h=-t+b\]

zepdrix (zepdrix):

To find our y-intercept (b), we can simply plug in one of the coordinate pairs and solve for b.

zepdrix (zepdrix):

How bout we useeeee, (5,15).

OpenStudy (anonymous):

so we're doing h = -5 + 15? kinda confused.

zepdrix (zepdrix):

They gave us these "coordinate pairs" to plug in for the `h` and `t` values in the equation.

zepdrix (zepdrix):

\[\Large (\color{orangered}{t},\;\color{royalblue}{h})=(\color{orangered}{5},\color{royalblue}{15})\]We want to plug these values \[\Large \color{royalblue}{h}=-\color{orangered}{t}+b\]into our equation:

zepdrix (zepdrix):

Blah that pasted a little funky lol

OpenStudy (anonymous):

Lol, well soo... 15 = -5 + b and we're trying to figure out b.

zepdrix (zepdrix):

Yah that sounds right.

OpenStudy (anonymous):

So then what? do you do the thing where you switch b with 15 (basically) but + changes to -?

zepdrix (zepdrix):

Hmmm, no. You're over thinking this I guess. :( 15 = -5 + b T solve for b, add 5 to each side.

OpenStudy (anonymous):

So doing so, it would be 20 = b since -5 cancels out...?

OpenStudy (anonymous):

b = 20??

zepdrix (zepdrix):

yup!

OpenStudy (anonymous):

YESS!!

zepdrix (zepdrix):

So that finishes the first part that we needed to do. `Finding a linear equation for our relationship`.\[\Large h=-t+20\]

zepdrix (zepdrix):

Do you understand what the 20 represents? That is the initial h value. So when t = 0, h = 20. Hmm, what does that mean?

OpenStudy (anonymous):

20 = 0 + 20...?

zepdrix (zepdrix):

No, i mean.. what does that number `represent` ?

OpenStudy (anonymous):

oh, lol XD it can either be time or height...and it has to be height because 20 represents the initial h value, right?

zepdrix (zepdrix):

Ya :o Initial height, right? So it's telling us that the candle was 20 inches long before it started burning.

zepdrix (zepdrix):

Just an interesting little tid bit there +_+ ya.

OpenStudy (anonymous):

It helps though! :D

zepdrix (zepdrix):

Now that you've got your linear equation set up, they want you to find `h` when `t=8`.

OpenStudy (anonymous):

h = -8 + 20?

OpenStudy (anonymous):

(which is 12 btw)

zepdrix (zepdrix):

yay team \c:/ looks good.

OpenStudy (anonymous):

That's it?? :D

zepdrix (zepdrix):

mhm

OpenStudy (anonymous):

Thank you so much! I thought I was dead in water there for a moment :DDDDD